A rectangle is 4 times as long as it is wide. A second rectangle is 5 centimenters longer and 2 centimeters wider than the first. The area of the second rectangle is 270 square centimeters greater than the first. What are the deminsions of the original rectangle?
Thank you in advance!!

Answer by scott8148(6628)

let x=width, so 4x=length

(x+2)(4x+5)=(x)(4x)+270 __ 4x^2+13x+10=4x^2+270

13x=260